OpenStudy (anonymous):

A kite is currently flying at an altitude of 14 meters above the ground. If the angle of elevation from the ground to the kite is 68°, find the length of the kite string to the nearest meter

OpenStudy (anonymous):

Sin 68 = 14/hyp hyp= 14/(sin68) =15.1m to 1 dp